OpenStudy (anonymous):

simplify 2a+2 over a^2-1

OpenStudy (anonymous):

the denominator is a diff of squares

OpenStudy (anonymous):

2 ----- (a - 1)?

OpenStudy (anonymous):

\(a^2-1=(a+1)(a-1)\) \(\dfrac{2a+2}{a^2-1}=\dfrac{2(a+1)}{(a+1)(a-1)}=\dfrac{2\cancel{(a+1)}}{\cancel{(a+1)}(a-1)}=\dfrac{2}{a-1}\)

OpenStudy (anonymous):

looks good

OpenStudy (anonymous):

thanks can you help me with some more questions?

OpenStudy (anonymous):

nah, sorry. trying to focus on some other stuff I have to do

OpenStudy (anonymous):

ok